OpenStudy (anonymous):

simplify 2x-24 over 8

OpenStudy (diyadiya):

x-12/4

OpenStudy (turingtest):

(2x-24)/8=2x/8-24/8=x/4-3

OpenStudy (turingtest):

Please put parentheses and not write the word "over" that's why you have two different answers...

OpenStudy (anonymous):

X-12 over 4